### v1

@@ 3.2 Retrieval and Generation Framework
Text Semantic Clustering. To efficiently retrieve multiple semantic relevant texts for each image, we initially encode all sentences using the EVA02-CLIP E/14-plus Sun et al. (2023) model. Inspired by Unicom An et al. (2023), we utilize the standard $K$-Means algorithm Ikotun et al. (2023) offline to partition the 0.84B texts into 2M clusters with the help of efficient feature quantization Johnson et al. (2019).
@@
Hierarchical Retrieval. Given the prohibitive computational overhead of direct semantic text retrieval from 0.84B sentences (exceeding 10,000 hours on 8 × A100 GPUs), we devise a hierarchical retrieval method to optimize computational efficiency. We first perform inter-cluster retrieval to find the most relevant cluster center for each image. Then, we group images sharing the same cluster center and perform intra-cluster retrieval to obtain multiple semantically relevant sentences. This approach enables the retrieval of 198M images and 0.84B sentences within 40 hours using 8 $\times$ A100 GPUs.
@@
Image Semantic Augmented Generation. Although the retrieved realistic texts achieve satisfactory performance, they exhibit limitations in capturing fine-grained visual semantics. To address this issue, we introduce the Image Semantic Augmented Generation module. This module initially employs the OFA model to generate a concise caption for each image. We then integrate the open-set image tagging model RAM++, which extracts object detection tags. Considering that RAM++ supports only 4,000 tags, we expand this set to 8,000 by incorporating an additional 4,000 tags derived from real-world sentences. Following CapsFusion Yu et al. (2024), we utilize ChatGPT4 Turbo to merge the retrieved realistic texts with concise captions and image tags to construct a 100K instruction dataset (the prompt we used is presented in the supplementary material). Subsequently, we fine-tune the LLaMA3-8B model Dubey et al. (2024) using the LLaMA Factory Zheng et al. (2024b) and deploy the vLLM Kwon et al. (2023) for large-scale inference. Ultimately, we convert data from 118M multimodal interleaved documents into 198M image-text pairs, where each image is associated with multiple retrieved realistic texts and synthetic texts.
3.3 Semantic Balance Sampling
To further improve the quality and diversity of our dataset, we implement semantic balancing sampling across the 198M image-text pairs. Specifically, we use EVA02-CLIP E/14-plus Sun et al. (2023) to encode and calculate the cosine similarity between imaged and synthetic texts. To reduce the impact of OCR-related or mismatched pairs during pre-training, we filter out 29.7M pairs with cosine similarities above 0.61 or below 0.51. Inspired by MetaCLIP Hu Xu (2023), we introduce an easy but efficient cluster-based semantic balance sampling strategy. We cluster the image embeddings from the remaining 168.3M pairs into 1M centers. To enhance the semantic diversity of our dataset, we randomly select 20, 35, and 180 samples from clusters exceeding these thresholds, while retaining all samples from smaller clusters. This approach culminates in the construction of the RealSyn15M, RealSyn30M, and RealSyn100M datasets.
4 Experiments and Results
4.1 Implementation Details
We initially collect 118M interleaved image-text documents from the OBELICS Laurençon et al. (2024) as our primary data source. We use OFAbase Wang et al. (2022) and RAM++large Huang et al. (2023) to generate brief captions and semantic tags. To validate the dataset performance, we pre-train standard CLIP supervised by the text randomly selected from the three retrieved realistic texts and one synthetic text. During pre-training, we adopt AdamW Loshchilov (2019) as the optimizer, with a learning rate of 1e-3 and a weight decay of 0.2. The parameters $\beta 1$ and $\beta 2$ are set to 0.9 and 0.98, respectively. The input image size is 224×224, and the input text sequence length is 77. The temperature parameter $\tau$ is initialized to 0.07. We train 32 epochs with 4096 batch sizes on 8 $\times$ A100 (80G) GPUs. Please refer to the supplementary material for more details.
To validate the effectiveness of the RealSyn dataset, we compare RealSyn with the previous datasets across various models and data scales. We compare RealSyn15M with the YFCC15M filtered by DeCLIP Li et al. (2022). Following ALIP Yang et al. (2023), we also compare with LAION15M, LAION30M, and LAION100M (subset randomly selected from LAION400M).
@@ 4.1 Implementation Details
Table 3: Zero-shot image-text retrieval performance on Flickr30k and MSCOCO. Pre-training CLIP-B/32 on RealSyn dataset achieves a significant improvement on all metrics.
4.2 Main Results
Linear Probe. In Tab. 1, we present the linear probe performance of the ViT-B/32 model across 20 downstream datasets. When pretrained on the 15M scale, RealSyn15M surpasses YFCC15M in 16 out of 20 datasets, registering an average performance increase of 6.9%. Moreover, RealSyn15M outperforms LAION15M in 18 out of 20 datasets, with an average improvement of 1.6%. Upon scaling the dataset to 30M and 100M, RealSyn achieves average performance improvement of 1.3% and 1.4% over LAION, respectively. These results demonstrate the efficacy of the RealSyn dataset in vision-language representation learning.
Zero-shot Transfer. We evaluate the zero-shot transfer performance of the ViT-B/32 model across 20 classification benchmarks using the same prompt templates as SLIP Mu et al. (2022). As indicated in Tab. 2, RealSyn15M surpasses YFCC15M on 18 of 20 datasets, achieving an average performance improvement of 14.3%. In comparison to LAION15M, RealSyn15M excels on 18 of 20 datasets with an average improvement of 5.2%. Upon expanding the dataset sizes to 30M and 100M, RealSyn achieves average performance improvements of 3.5% and 2.3% compared to LAION, highlighting its efficiency and scalability.
It is important to note that RealSyn demonstrates a significant decrease in performance on certain datasets, such as Cars and Flowers. This reduction is primarily attributed to the unique data distribution of RealSyn, characterized by a scarcity of data for specific concepts, which hampers the model’s ability to effectively learn these concepts. For instance, as depicted in Fig. 4, samples related to cars constitute merely 0.9% of the dataset.
@@ 4.2 Main Results
Table 4: Zero-shot robustness comparison. Pre-training CLIP-B/32 on RealSyn demonstrates superior robustness across all datasets.
Zero-shot Image-Text Retrieval. In Tab. 3, we present the zero-shot image-text retrieval performance of the ViT-B/32 model pre-trained on different scale of datasets. RealSyn achieves superior results across all evaluation metrics. Specifically, RealSyn15M improves Recall@1 by 35.8%&26% on Flickr30K Young et al. (2014) and by 22.5%&12.6% on MSCOCO Lin et al. (2014). RealSyn30M improves Recall@1 by 16.4%&11.6% on Flickr30K Young et al. (2014) and by 12.3%&7.4% on MSCOCO Lin et al. (2014). This significant enhancement in cross-modal retrieval performance indicates that the RealSyn dataset effectively improves vision-language representation learning by utilizing realistic and synthetic texts, resulting in robust representations and enhanced cross-modal alignment.
@@
Figure 4: A T-SNE Van der Maaten and Hinton (2008) projection of LDA Blei et al. (2003) topic cluster from a randomly selected 1M samples from RealSyn.
Zero-shot Robustness. In Tab. 4, we present the zero-shot robustness performance. The results show that RealSyn significantly improves the robustness of vision-language pre-training models. Specifically, compared to LAION, RealSyn increases average performance by 4.3%, 4.2%, and 2.8% for datasets of 15M, 30M, and 100M, respectively. This notable improvement in performance primarily stems from the use of retrieved realistic texts that are not bound by the limitations of generative models, coupled with superior conceptual diversity compared to YFCC and LAION, thereby substantially enhancing model robustness.
5 Analysis
@@
Figure 5: The richness assessment and diversity assessment on different datasets. RealSyn-R1: the most relevant retrieved realistic text. RealSyn-S1: the semantic augmented synthetic text based on RealSyn-R1.
5.1 Statistics Analysis
@@
Topic-based Assessment. Following MMC4 Zhu et al. (2024), we ran LDA Blei et al. (2003) on random sampling 1M image-realistic text pairs with 30 topics. Fig. 4 presents the proportions and examples for six topics: animal, food, airplane, flower, automotive, and landmark. Notably, the dataset contains minimal samples related to “flower” and “automotive” topics, representing merely 0.4% and 0.9% of the total, respectively. This paucity of examples hinders the model’s ability to sufficiently learn these concepts, thereby compromising its performance in the linear probe and zero-shot transfer evaluations on the Flowers and Cars datasets.
@@
Richness Assessment. Fig. 5(a) presents image-text similarity and text token distribution of 15M samples from YFCC15, LAION, RealSyn-R1 (the most relevant retrieved realistic text), and RealSyn-S1 (the semantic augmented synthetic text based on RealSyn-R1). Compared to the datasets collected from the Internet, RealSyn exhibits robust similarity metrics, even after removing OCR data. Moreover, both the retrieved realistic texts and synthetic texts contain a larger quantity of words, which can provide a richer textual context that enhances vision-language representation learning.
@@
Diversity Assessment. The RealSyn is constructed based on real-world interleaved image-text documents, which encompasses a wide array of diverse information. Following previous work Lai et al. (2025), we randomly select 0.2M samples to calculate the number of unique entities in the caption to assess the data diversity of different datasets. As depicted in Fig. 5(b), both the retrieved realistic texts and image semantic augmented synthetic texts exhibit a higher number of distinct entities. Such diversity enriches the dataset, facilitating the model’s acquisition of comprehensive knowledge and enhancing both performance and robustness.
@@
Model Scaling. In Sec. 4.2, RealSyn exhibits superior performance across various data scales (More data scaling analysis is presented in the supplementary material). To further explore the model scaling capability, we present the downstream task performance of three models in Fig. 6. Notably, compared to LAION, RealSyn demonstrates steeper slopes in performance curves across linear probing, zero-shot transfer, and robustness, indicative of its superior model scaling capabilities.
@@
Figure 6: Model scaling capability. We compare the models pre-trained on LAION30M and RealSyn30M.
Figure 7: Image captioning comparisons on COCO2017 and Flickr30k. B4, MT., RL. and Cd. represent the metric of BLEU, METEOR, ROUGE-L, and Cider.
5.2 Ablation Study
Image Captioning via MLLM. Following LLaVA-1.5 Liu et al. (2024), we initially align visual features to the text space using 558k data. Subsequently, we construct an image captioning dataset from LAION and RealSyn, which we then use for instruction tuning. Specifically, we split the realistic text and synthetic text for the same image into two independent samples, totaling 2M samples for training for one epoch. Concurrently, we randomly select 1M samples from LAION and train for two epochs. As illustrated in Fig. 7, RealSyn demonstrates a significant enhancement in performance compared to LAION across all evaluation metrics on both the COCO2017 Lin et al. (2014) and Flickr30k Young et al. (2014).
Extension to Pure Image. To further extend our method for pure images, we conduct experiments on ImageNet Deng et al. (2009). Initially, we retrieve semantically relevant realistic texts for each ImageNet image from our sentence database and generate image semantic augmented synthetic texts. Then, we pre-train ResNet50 He et al. (2016) supervised by the text randomly selected from the retrieved realistic texts and synthetic texts. Comparative analysis with SimCLR Chen et al. (2020) under identical conditions shows a linear probe average performance enhancement of 2.1% across 12 datasets using our constructed data. Detailed experimental results are provided in the supplementary material.
@@ 5.2 Ablation Study
Table 5: Comparison of concept balance sampling and random sampling on the 15M dataset.
Ablation on Semantic Balance Sampling. To demonstrate the effectiveness of our proposed semantic balance sampling method, we compare it with random sampling. As indicated in Tab. 5, concept balance sampling yields performance improvements of 0.7%, 1.1%, and 1.0% in linear probe, zero-shot transfer, and robustness, respectively. Besides, we visualize the data distribution using different sampling methods by clustering 15M samples into 1M centers. As shown in Fig. 8, the distribution from semantic balanced sampling is smoother, facilitating the learning of long-tail concepts.
Figure 8: Clustering distribution of 15M data obtained from random sampling and semantic balanced sampling.
@@
Table 6: Ablation experiment results using different types of text on the 15M dataset. $T^{k}_{r}$: the $k$-th retrieved semantic relevant realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$.

### v3

@@ 3.2. Retrieval and Generation Framework
Text Semantic Clustering. To efficiently retrieve multiple semantically relevant texts for each image, we initially encode all sentences using the EVA02-CLIP E/14-plus (Sun et al., 2023) model. Inspired by Unicom (An et al., 2023), we utilize the standard $K$-Means algorithm (Ikotun et al., 2023) offline cluster all sentences into a specified number of clusters. To minimize intra-cluster and inter-cluster conflicts, we initially establish the optimal number of cluster centroids through small-scale experiments. Subsequently, we divide the 0.84 billion texts into two million clusters, utilizing efficient feature quantization techniques (Johnson et al., 2019).
@@
Hierarchical Retrieval. Given the prohibitive computational cost of direct semantic text retrieval across 0.84B sentences (requiring over 10,000 GPU-hours on 8×A100), we propose a hierarchical retrieval framework to enhance efficiency. Given an image, we perform inter-cluster retrieval to identify the most relevant cluster center for the image. Subsequently, we conduct intra-cluster retrieval within the identified cluster to obtain multiple real-world sentences semantically matched to the image. This approach achieves scalable retrieval of 198M images and 0.84B sentences within 40 GPU-hours on an 8×A100.
@@
Image Semantic Augmented Generation. Although the retrieved realistic texts achieve satisfactory performance, they exhibit limitations in capturing fine-grained visual semantics. To address this issue, we introduce the Image Semantic Augmented Generation module. This module initially employs the OFA (Wang et al., 2022) model to generate a concise caption for each image. We then integrate the open-set image tagging model RAM++ (Huang et al., 2023), which extracts object detection tags. Considering that RAM++ supports only 4,000 tags, we expand this set to 8,000 by incorporating an additional 4,000 tags derived from real-world sentences. Following CapsFusion (Yu et al., 2024), we utilize ChatGPT4 Turbo to merge the retrieved realistic texts with concise captions and image tags to construct a 100K instruction dataset (the prompt we used is presented in the supplementary material). Subsequently, we conduct instruction tuning of the LLaMA3-8B model (Dubey et al., 2024) using the LLaMA Factory (Zheng et al., 2024c) and deploy the vLLM (Kwon et al., 2023) for large-scale inference. Ultimately, we convert data from 118M multimodal interleaved documents into 198M image-text pairs, where each image is associated with multiple retrieved realistic texts and synthetic texts.
Table 3. Zero-shot image-text retrieval performance on Flickr30k and MSCOCO. Pre-training CLIP-B/32 on RealSyn dataset achieves a significant improvement on all metrics.
@@
Table 4. Zero-shot robustness comparison on different IN-1K val set variants. Pre-training CLIP-B/32 on RealSyn demonstrates superior robustness across all datasets.
@@
3.3. Semantic Balance Sampling
To further improve the quality and diversity of our dataset, we implement semantic balancing sampling across the 198M image-text pairs. Specifically, we use EVA02-CLIP E/14-plus (Sun et al., 2023) to encode and calculate the cosine similarity between images and synthetic texts. To reduce the impact of OCR-related or mismatched pairs during pre-training, we filter out 29.7M pairs with cosine similarities scores either above 0.61 or below 0.51, as determined through human verification. Inspired by MetaCLIP (Hu Xu, 2023), we introduce an easy but efficient cluster-based semantic balance sampling strategy. We cluster the image embeddings from the remaining 168.3M pairs into 1M centers. To enhance the semantic diversity of our dataset, we randomly select 20, 35, and 180 samples from clusters exceeding these thresholds, while retaining all samples from smaller clusters. This approach culminates in the construction of the RealSyn15M, RealSyn30M, and RealSyn100M datasets.
4. Experiments and Results
4.1. Implementation Details
We initially collect 118M interleaved image-text documents from the OBELICS (Laurençon et al., 2024) as our primary data source. We use OFAbase (Wang et al., 2022) and RAM++large (Huang et al., 2023) to generate brief captions and semantic tags. To validate the dataset performance, we pre-train standard CLIP supervised by the text randomly selected from the three retrieved realistic texts and one synthetic text, inspired by the LaCLIP (Fan et al., 2024). During pre-training, we adopt AdamW (Loshchilov, 2019) as the optimizer, with a learning rate of 1e-3 and a weight decay of 0.2. The parameters $\beta 1$ and $\beta 2$ are set to 0.9 and 0.98, respectively. The input image size is 224×224, and the input text sequence length is 77. The temperature parameter $\tau$ is initialized to 0.07. We train 32 epochs with 4096 batch sizes on 8 $\times$ A100 (80G) GPUs. Please refer to the supplementary material for more details.
To validate the effectiveness of the RealSyn dataset, we compare RealSyn with the previous datasets across various models and data scales. We compare RealSyn15M with the YFCC15M filtered by DeCLIP (Li et al., 2022). Following ALIP (Yang et al., 2023), we also compare with LAION15M, LAION30M, and LAION100M (subset randomly selected from LAION400M).
@@ 4.1. Implementation Details
Figure 4. A T-SNE (Van der Maaten and Hinton, 2008) projection of LDA (Blei et al., 2003) topic cluster from a randomly selected 1M samples from RealSyn. RealSyn encompasses a broad range of everyday topics, e.g., animal, food, airplane, etc. We also display representative central images for each identified topic.
4.2. Main Results
Linear Probe. In Table 1, we report the linear probe performance of the ViT-B/32 model across 20 downstream datasets. When pretrained at the 15M scale, RealSyn15M exceeds YFCC15M on 16 of 20 datasets, achieving an average performance increase of 6.9%. Additionally, RealSyn15M outperforms LAION15M on 18 of 20 datasets, with an average improvement of 1.6%. With dataset scaling to 30M and 100M, RealSyn achieves average performance improvements of 1.3% and 1.4% over LAION, respectively. These findings underscore the superior CLIP training efficiency of the RealSyn dataset compared to the widely-used YFCC and LAION datasets across multiple scales.
Zero-shot Transfer. We evaluate the zero-shot transfer performance of the ViT-B/32 model across 20 classification benchmarks using the same prompt templates as SLIP (Mu et al., 2022). As indicated in Table 2, RealSyn15M surpasses YFCC15M on 18 of 20 datasets, achieving an average performance improvement of 14.3%. In comparison to LAION15M, RealSyn15M excels on 18 of 20 datasets with an average improvement of 5.2%. Upon expanding the dataset sizes to 30M and 100M, RealSyn achieves average performance improvements of 3.5% and 2.3% compared to LAION, highlighting its efficiency and scalability.
It is important to note that RealSyn demonstrates a significant decrease in performance on certain datasets, such as Cars and Flowers. This reduction is primarily attributed to the unique data distribution of RealSyn, characterized by a scarcity of data for specific concepts, which hampers the model’s ability to effectively learn these concepts. For example, as shown in Figure 4, samples related to cars and flowers represent only 0.9% and 0.4% of the dataset, respectively.
Zero-shot Image-Text Retrieval. In Table 3, we present the zero-shot image-text retrieval performance of the ViT-B/32 model pre-trained on different scales of datasets. RealSyn achieves superior results across all evaluation metrics. Specifically, RealSyn15M improves Recall@1 by 35.8%&26% on Flickr30K (Young et al., 2014) and by 22.5%&12.6% on MSCOCO (Lin et al., 2014). RealSyn30M improves Recall@1 by 16.4%&11.6% on Flickr30K (Young et al., 2014) and by 12.3%&7.4% on MSCOCO (Lin et al., 2014). RealSyn100M improves Recall@1 by on 14.6%&8.4% Flickr30K (Young et al., 2014) and by 9%&5.4% on MSCOCO (Lin et al., 2014). This significant enhancement in cross-modal retrieval performance indicates that the RealSyn dataset effectively improves contrastive vision-language representation learning by incorporating both realistic and synthetic texts, resulting in robust representations and enhanced cross-modal alignment.
Zero-shot Robustness. To further validate the robustness of the RealSyn dataset, we present the zero-shot robustness (Radford et al., 2021) performance across various IN-1K (Deng et al., 2009) val set variants in Table 4. The results indicate that RealSyn significantly enhances the robustness of vision-language pre-training models. Specifically, RealSyn15M shows a performance increase of 13.1% and 4.3% compared to YFCC15M and LAION15M, respectively. When scaled to 30M and 100M, RealSyn achieves additional improvements of 4.2% and 2.8% over LAION. This substantial enhancement in performance is primarily attributable to the utilization of retrieved realistic texts, which surpass the constraints of generative models, and the superior conceptual diversity relative to YFCC and LAION, thus significantly boosting model robustness.
Figure 5. Image captioning comparisons on COCO2017 and Flickr30k. B4, MT., RL. and Cd. represent the metric of BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Banerjee and Lavie, 2005), and Cider (Vedantam et al., 2015).
Image Captioning via MLLM. In Figure 5, we present the image captioning performance of the LLaVA-1.5 (Liu et al., 2024) trained using different datasets (LAION v.s. RealSyn). Initially, we first map visual features into the textual domain using the initial 558k dataset from LLaVA-1.5. We then develop an image captioning dataset from both LAION and RealSyn for instruction tuning. Specifically, we select 1M samples randomly from each dataset and train over two epochs. As depicted in Figure 5, RealSyn significantly outperforms LAION in all evaluation metrics on both the COCO2017 (Lin et al., 2014) and Flickr30k (Young et al., 2014) benchmarks. This notable performance enhancement confirms the higher quality and better image-text alignment of the RealSyn dataset.
@@ 4.2. Main Results
Figure 6. The richness assessment and diversity assessment on different datasets. RealSyn-R1: the most relevant retrieved realistic text. RealSyn-S1: the semantic augmented synthetic text based on RealSyn-R1.
5. Analysis
Figure 7. Data Scaling Analysis. We pretrain ViT-B/32 on RealSyn in different data scales. $\times$ represents the results predicted using our data scaling law.
5.1. Statistics Analysis
@@
Topic-based Assessment. Following MMC4 (Zhu et al., 2024b), we ran LDA (Blei et al., 2003) on random sampling 1M image-realistic text pairs with 30 topics. Figure 4 presents the proportions and examples for six topics: animal, food, airplane, flower, automotive, and landmark. Notably, the dataset contains minimal samples related to “flower” and “automotive” topics, representing merely 0.4% and 0.9% of the total, respectively. This paucity of examples hinders the model’s ability to sufficiently learn these concepts, thereby compromising its performance in the linear probe and zero-shot transfer evaluations on the Flowers and Cars datasets.
@@
Richness Assessment. Figure 6(a) presents image-text similarity and text token distribution of 15M samples from YFCC15, LAION, RealSyn-R1 (the most relevant retrieved realistic text), and RealSyn-S1 (the semantic augmented synthetic text based on RealSyn-R1). Compared to datasets sourced from the Internet, RealSyn demonstrates robust similarity metrics even after the removal of OCR data. This effectively explains the significant performance enhancement observed in the CLIP model trained on RealSyn for image-text retrieval tasks. Moreover, both the retrieved realistic texts and synthetic texts contain a larger quantity of words, which can provide a richer textual context that enhances vision-language representation learning.
@@
Diversity Assessment. The RealSyn is constructed based on real-world interleaved image-text documents, which encompasses a wide array of diverse information. Following previous work (Lai et al., 2025), we randomly select 0.2M samples to calculate the number of unique entities in the caption to assess the data diversity of different datasets. As depicted in Figure 6(b), both the retrieved realistic texts and image semantic augmented synthetic texts exhibit a higher number of distinct entities. Such diversity enriches the dataset, facilitating the model’s acquisition of comprehensive knowledge and enhancing both performance and robustness.
Data Scaling Analysis. We present the data scaling law (Kaplan et al., 2020) derived from our RealSyn dataset, justifying its scalability over samples. Specifically, we conduct a series of visual-language pre-trainings with proposed datasets ranging from 12M to 60M, and fit each performance metric to the inverse of logarithmic functions with respect to the number of millions of training samples $x$. Based on the fitting results from these preliminary experiments, we extrapolate each performance scaling law to $100M$ samples, and validate their predicted scaling trends with our RealSyn100M100M dataset as shown in Figure 7. Notably, as indicated by the coefficients shown in Eq. 3, these performance laws also likely suggest an upper bound of model capability that a ViT-B/32 could possibly reach through our proposed visual-language pre-training paradigm with multimodal interleaved documents:
(3) $\text{Linear Probe:}\ \mathcal{L}(x)$ $0\displaystyle\approx\frac{-0.21}{log(x-4.23)}+8$0
[…]
@@
Model Scaling Analysis. In Section 4.2, RealSyn exhibits superior performance across various data scales. To further explore the model scaling capability, we present the downstream task performance of three models in Figure 8. Notably, compared to LAION, RealSyn demonstrates steeper slopes in performance curves across linear probe, zero-shot transfer, and robustness, indicative of its superior model scaling capabilities.
@@
Figure 8. Model scaling capability. We compare the models pre-trained on LAION30M and RealSyn30M.
Table 5. Comparison of semantic balance sampling and random sampling on the 15M dataset.
@@
Table 6. Ablation study examining the quantity and combinations of retrieved realistic texts and corresponding image semantic augmented synthetic texts within a 15M-scale dataset. $T^{k}_{r}$: the $k$-th retrieved semantic relevant realistic text. $T^{k}_{s}$: the image semantic augmented synthetic text for $T^{k}_{r}$.
@@
$T^{1}_{r}$ $T^{2}_{r}$ $T^{3}_{r}$ $T^{1}_{s}$ Linear probe Avg Transfer Avg Robustness Avg
70.3 42.4 25.7
71.2 46.8 30.7
70.2 39.7 24.0
71.4 47.9 31.5
Extension to Pure Image. To further extend our transformation paradigm for pure images, we conduct experiments on ImageNet (Deng et al., 2009). Initially, we retrieve semantically relevant realistic texts for each ImageNet image from our sentence database and generate image semantic augmented synthetic texts. Then, we pre-train ResNet50 (He et al., 2016) supervised by the text randomly selected from the retrieved realistic texts and synthetic texts. Comparative analysis with SimCLR (Chen et al., 2020) under identical conditions shows a linear probe average performance enhancement of 2.1% across 12 datasets using our constructed data. Detailed experimental results are provided in the supplementary material.
5.2. Ablation Study
Ablation on Semantic Balance Sampling. To demonstrate the efficacy of our proposed semantic balance sampling method, we contrast it with random sampling. As shown in Table 5, semantic balance sampling achieves performance improvements of 0.7%, 1.1%, and 1.0% in linear probe accuracy, zero-shot transfer, and robustness, respectively. Besides, we visualize the data distribution of 15M samples clustered into 1M centers using different sampling strategies. As shown in Figure 9 reveals that semantic balance sampling results in a smoother distribution, thereby enhancing the learning of long-tail concepts.
Figure 9. Clustering distribution of 15M data obtained from random sampling and semantic balance sampling.